### v3

# Integer quadratic forms and extensions of subsets of linearly independent roots

Abstract.
We consider subsets of linearly independent roots in a certain root system $\varPhi$. Let $S^{\prime}$ be such a subset, and let $S^{\prime}$ be associated with any Carter diagram $\Gamma^{\prime}$. The main question of the paper: what root $\gamma\in\varPhi$ can be added to $S^{\prime}$ so that $S^{\prime}\cup\gamma$ is also a subset of linearly independent roots? This extra root $\gamma$ is called the linkage root. The vector $\gamma^{\nabla}$ of inner products $\{(\gamma,\tau^{\prime}_{i})\mid\tau^{\prime}_{i}\in S^{\prime}\}$ is called the linkage label vector. Let $B_{\Gamma^{\prime}}$ be the Cartan matrix associated with $\Gamma^{\prime}$. It is shown that $\gamma$ is a linkage root if and only if $\mathscr{B}^{\vee}_{\Gamma^{\prime}}(\gamma^{\nabla})<2$, where $\mathscr{B}^{\vee}_{\Gamma^{\prime}}$ is a quadratic form with the matrix inverse to $B_{\Gamma^{\prime}}$. The set of all linkage roots for $\Gamma^{\prime}$ is called a linkage system and is denoted by $\mathscr{L}(\Gamma^{\prime})$. The Cartan matrix associated with any Carter diagram $\Gamma^{\prime}$ is conjugate to the Cartan matrix associated with some Dynkin diagram $\Gamma$ [St23]. The sizes of $\mathscr{L}(\Gamma^{\prime})$ and $\mathscr{L}(\Gamma)$ are the same. Let $W^{\vee}$ be the Weyl group of the quadratic form $\mathscr{B}^{\vee}_{\Gamma^{\prime}}$. This group acts on the linkage system and forms several orbits. The sizes and structure of orbits for linkage systems $\mathscr{L}(D_{l})$ and $\mathscr{L}(D_{l}(a_{k}))$ are presented.
1. Introduction
1.1. Brief summary
[…]
Let $\widetilde{\Gamma}$ be the diagram obtained from a Carter diagram $\Gamma$ by adding one extra root $\gamma$, with its bonds, so that the roots corresponding to $\widetilde{\Gamma}$ form a linearly independent root subset. We would like to describe such roots in the general case. The extra root $\gamma$ is called a linkage root, the diagram $\widetilde{\Gamma}$ is called a linkage diagram. More precise definitions will be given shortly.
The simple criterion for a given vector $\gamma$ to be a linkage root given in the terms of the inverse quadratic form $\mathscr{B}^{\vee}$ to the Cartan matrix associated with $\Gamma$ is as follows:
$\mathscr{B}^{\vee}_{\Gamma}(\gamma^{\nabla})<2,$ (1.1)
where $\gamma^{\nabla}$ is so-called linkage label vector, the vector of the inner products of $\gamma$ with roots from a given $\Gamma$-set (Theorem 3.3). There exists one-to-one correspondence between linkage diagrams and linkage label vectors. The Weyl group $W^{\vee}$ preserving the quadratic form $\mathscr{B}^{\vee}$ acts on the space of the linkage diagrams, or, equivalently, on the space of linkage label vectors, see §2.4.
The set $\mathscr{L}_{\widetilde{\Gamma}}(\Gamma)$ of all linkage systems for the given vertex extension $\Gamma\prec\widetilde{\Gamma}$ is said to be partial linkage system. (For the definition of vertex extension, see §4.1.) The union of all partial systems using all possible extensions $\Gamma\prec\widetilde{\Gamma}$ is called the full linkage system of $\Gamma$, or, for short, linkage system:
$\mathscr{L}(\Gamma^{\prime})=\bigcup_{\Gamma^{\prime}\prec\widetilde{\Gamma}}\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime}).$ (1.2)
The size and structure of linkage systems $\mathscr{L}(D_{l})$ and $\mathscr{L}(D_{l}(a_{k}))$ are described in Theorem 5.3, and Figs. 912, Figs. 45, see also Table 2.1. The group $W^{\vee}$ acts on the linkage system and forms several orbits that are not connected to each other.
The Carter diagrams having the same rank and type form homogeneous classes. For example, $E_{6}$, $E_{6}(a_{1})$ and $E_{6}(a_{2})$ form the homogeneous class $C(E_{6})$. Any homogeneous class contains only one Dynkin diagram, see §1.5. For each pair of diagrams $\{\widetilde{\Gamma},\Gamma\}$ out of one homogeneous class, there exists transition matrix $M$ mapping each $\widetilde{\Gamma}-$set $\widetilde{S}$ to some $\Gamma$-set $S$ (Theorem 1.1). The matrix $M$ acts only on one element in $\widetilde{S}$, the remaining elements remain fixed. The transition matrix is an involution. Transition matrices make it possible to connect objects associated with different Carter diagrams from the same homogeneous class, see §1.4
The transition mappings are close in properties to the transformations introduced by A.Gabrielov for systematic study of quadratic forms associated with singularities, and to the transformations introduced by S.Ovsienko for the study weakly positive unit quadratic forms, see §1.6.1.
1.2. Root system, Cartan matrix, Carter diagram
Let $\varPhi$ be a finite root system, $W$ be the corresponding finite Weyl group. We consider only root systems with simply-laced Dynkin diagrams. The Cartan matrix associated with $\varPhi$ is denoted by $B$. Let $\mathscr{B}$ be the quadratic Tits form associated with $B$ and $(\cdot,\cdot)$ be the inner product induced by $\mathscr{B}$. For the Cartan matrix $B$, the following well-known property [^1] For the matrix $B/2$, eq. (1.3) looks like this: $\mathscr{B}(\alpha)=1\Longleftrightarrow\alpha\in\varPhi$, see [K80, Prop. 1.6]. holds:
$\mathscr{B}(\alpha)=2\Longleftrightarrow\alpha\in\varPhi.$ (1.3)
In [Ca72], Carter introduced admissible diagrams to describe conjugacy classes in $W$. These diagrams can be used to describe other objects, such as root subsets [^2] In what follows, the phrase “root subset” means a subset of linearly independent roots. and quadratic forms. Let $S\subset\varPhi$ be a root subset, roots of $S$ are not necessarily simple. To the subset $S$ we associate some diagram $\Gamma$ that provides one-to-one correspondence between roots of $S$ and nodes of $\Gamma$. The diagram $\Gamma$ is said to be admissible if the following two conditions hold:
[…]
In admissible diagrams all edges are solid, whereas Carter diagrams have solid and dotted diagrams that distinguish negative and positive inner products, see §1.7. In [St17], it was observed that the cycles in the Carter diagrams contain at least one solid edge $\{\alpha_{1},\beta_{1}\}$ (on which $(\alpha_{1},\beta_{1})<0$) and at least one dotted edge $\{\alpha_{2},\beta_{2}\}$~(on which $(\alpha_{2},\beta_{2})>0$), see §1.9. In this paper we introduce new objects - linkage roots, linkage diagrams and others, which are also closely related to Carter diagrams.
1.3. The partial Cartan matrix
Let $\Gamma$ be a Carter diagram. A set of linearly independent roots $S=\{\tau_{1},\dots,\tau_{n}\}$ is called $\Gamma$-set, if the vertices of the diagram $\Gamma$ are on one-to-one correspondence with the roots of $S$ and the inner products $(\tau_{i},\tau_{j})$ correspond to the edges of $\Gamma$ as follows:
[…]
Besides this, $(\tau_{i},\tau_{i})=2$ for any $i$. Similarly to the Cartan matrix associated with Dynkin diagrams, for each pair $\Gamma$, S}, we determine the following matrix:
$B_{\Gamma}:=\left(\begin{array}[]{cccccc}(\tau_{1},\tau_{1})&\dots&(\tau_{1},\tau_{n})\\ \dots&\dots&\dots\\ (\tau_{n},\tau_{1})&\dots&(\tau_{n},\tau_{n})\\ \end{array}\right).$ (1.5)
We call the matrix $B_{\Gamma}$ the partial Cartan matrix corresponding to $\Gamma$. Why partial? Since in this case some off-diagonal elements are $1$, they correspond to the dotted edges. The partial Cartan matrix $B_{\Gamma}$ is well-defined since products $(\tau_{i},\tau_{j})$ in (1.5) do not depend on the choice of the $\Gamma$-set $S$. The elements of the partial Cartan matrix are uniquely determined by the diagram $\Gamma$.
1.4. Some properties of Dynkin and Carter diagrams
Any simply-laced Dynkin diagram is the particular case of the Carter diagram. We will use the phrase “Carter diagram” to refer to a Carter diagram which is not a Dynkin diagram unless otherwise stated. For a list of all Carter diagrams, see Fig. 1. This list is divided into homogeneous classes, see §1.5.
(i) Consider the orbit of any root subset associated with a Carter diagram (resp. Dynkin diagram) under the action of the Weyl group. In the case of a Carter diagram, each root subset in this orbit contains non-simple roots. In the case of a Dynkin diagram, some root subset on this orbit consists only of simple roots.
(ii) For both Carter and Dynkin diagrams, the corresponding quadratic forms are positive definite; for both of them the spectrum of the corresponding Cartan matrix lies in the open interval $(0,4)$, see Propositions 3.1 and 3.5. The latter follows from the fact that the partial Cartan matrix $B_{\Gamma^{\prime}}$ for the Carter diagram $\Gamma^{\prime}$ is related to the Cartan matrix $B_{\Gamma}$ for some Dynkin diagram $\Gamma$ by a certain transition matrix $M$:
$^{t}{M}\cdot B_{\Gamma^{\prime}}\cdot M=B_{\Gamma},$ (1.6)
see §1.6.
(iii) The transition matrix $M$ modifies the root subset, the quadratic form as (1.6) and even the underlying diagram. The transition matrix combines the Carter diagrams and the corresponding Dynkin diagram together into one homogeneous class, see §1.5. Each homogeneous class contains only one Dynkin diagram.
(iv) The Carter diagrams contain $4$-cycles $D_{4}(a_{1})$ as subdiagrams. In principle, Carter diagrams can contain the cycles of length > $4$. It is shown in [St17, Th 3.1] that any Carter diagram containing $l$-cycles, where $l>4$, is equivalent to another Carter diagram containing only $4$-cycles. To realize this equivalence, an explicit transformation was constructed that maps each Carter diagram with long cycles into some Carter diagram containing only $4$-cycles.
(v) The correspondence between root subsets and Carter diagrams is not one-to-one. For example, the set of four linear independent roots $\{\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}\}$ connected as shown in Fig. 5, can be associated with exactly two conjugacy classes having the following representatives: $w_{o}=s_{\alpha_{1}}s_{\beta_{1}}s_{\alpha_{2}}s_{\beta_{2}}$ (associated with $D_{4}$) and $w_{1}=s_{\alpha_{1}}s_{\alpha_{2}}s_{\beta_{1}}s_{\beta_{2}}$ (associated with $D_{4}(a_{1})$), [St17, §§1.2.3–1.2.4]. The element $w_{o}$ corresponds to the Coxeter transformation class, and $w_{1}$ corresponds to the non-Coxeter class. As Vavilov and Migrin noted in [VM21, §1.2], this single phenomenon explains all non-Coxeter classes for all Carter diagrams with cycles.
Figure 1. Homogeneous classes of Carter diagrams $C(D_{l}),l\geq 4$ and $C(E_{l}),l=6,7,8$.
1.5. Homogeneous Carter diagrams and the transition matrix
The Dynkin diagrams $A_{l}$, where $l\geq 1$ (resp. $D_{l}$, $l\geq 4$; resp. $E_{l}$, $l=6,7,8$) are said to be the Dynkin diagrams of $A$-type (resp. $D$-type, resp. $E$-type). The Carter diagrams $D_{l}$, $D_{l}(a_{k})$, $l\geq 4$, $1\leq k\leq\big{[}\tfrac{l-2}{2}\big{]}$ (resp. $E_{l}$, $E_{l}(a_{k})$, $l=6,7,8$, $k$ are given in (1.7)) are said to be the Carter diagrams of $D$-type, resp. $E$-type).
The Carter diagrams of the same type and the same index constitute a homogeneous class of Carter diagrams. Denote by $C(\Gamma)$ the homogeneous class containing the Carter diagram $\Gamma$, see (1.7) and Fig. 1.
$\begin{array}[]{ll}&C(E_{6})=\{E_{6},E_{6}(a_{k}),k=1,2\},\\ &C(E_{7})=\{E_{7},E_{7}(a_{k}),1\leq k\leq 4\},\\ &C(E_{8})=\{E_{8},E_{8}(a_{k}),1\leq k\leq 8\},\\ &C(D_{l})=\{D_{l},D_{l}(a_{k}),1\leq k\leq\big{[}\tfrac{l-2}{2}\big{]}\},\text{ where }l\geq 4.\end{array}$ (1.7)
Let $\widetilde{S}$ (resp. $S$) be a $\widetilde{\Gamma}-$set (resp. $\Gamma$-set). In [St23, Tables 2-5], the transition matrices $M:\widetilde{S}\longmapsto S$ are constructed for the following homogeneous pairs $\{\widetilde{\Gamma},\Gamma\}$:
[…]
To our considerations, let us add diagram similarity mapping $L_{\tau_{i}}$ reflecting the root $\tau_{i}\in\widetilde{S}$ and fixing the remaining roots in $\widetilde{\Gamma}-$set:
$L_{\tau_{i}}:\tau_{i}\longmapsto-\tau_{i}.$ (1.9)
Eq. (1.9) means that any solid (resp. dotted) edge with vertex $\tau_{i}$ is mapped to the dotted (resp. solid) edge. Two Carter diagrams obtained from each other by a sequence of reflections (1.9), are called similar Carter diagrams. The diagrams obtained as images of the transition matrices $M$ in (1.6) are considered up to similarity of Carter diagrams, see [St17, §1.3]. The list (1.8) is called the adjacency list.
1.6. Transitions
The transition matrices proposed in [St23] transform root subsets associated with the corresponding Carter diagrams. Given a Carter diagram $\Gamma$, each transition changes only one root in the subset associated with $\Gamma$.
Theorem 1.1 (Transitions theorem).
[St23, Th. 4.1] For each pair of diagrams $\{\widetilde{\Gamma},\Gamma\}$ out of the list (1.8), there exists the transition matrix $M$ mapping each $\widetilde{\Gamma}-$set $\widetilde{S}$ to some $\Gamma$-set $S$. The matrix $M$ acts only on one element $\widetilde{\alpha}\in\widetilde{S}$ and does not change remaining elements in $\widetilde{S}$:
[…]
The image $S=M\widetilde{S}$ the $\Gamma$-set. The transition matrix $M:\widetilde{S}\longmapsto S$ is an involution:
[…]
The transition matrix change the underlying diagram for given subset of roots. For more details about the transitions matrices $M$ corresponding to the adjacency list (1.8), see Tables 2-5 in [St23, §4.6]. In §3 we use transitions $M$ to relate partial Cartan matrices associated with homogeneous Carter diagrams, as well as to relate linkage systems associated with them, see §1.8.
1.6.1. Transition mappings in other contexts
I would like to highlight two areas that use transformations close to the transition matrices from Theorem 1.1. In 1973, A. Gabrielov introduced transformations for the systematic study of quadratic forms associated with singularities, [G73], see Gabrielov’s examples in §A.1. In 1978, S. Ovsienko introduced so-called inflation technique for study the weakly positive unit quadratic forms, [O78]. For the Ovsienko theorem, see §A.2.
1.7. Admissible diagrams and Carter diagrams
Each element $w\in W$ can be expressed in the form
$w=s_{\alpha_{1}}s_{\alpha_{2}}\dots s_{\alpha_{k}},\text{ where }\alpha_{i}\in\varPhi\text{ for all }i.$ (1.11)
Carter proved that $k$ in the decomposition (1.11) is the smallest if and only if the subset of roots $\{\alpha_{1},\alpha_{2},\dots,\alpha_{k}\}$ is linearly independent; such a decomposition is said to be reduced. The admissible diagram corresponding to the given element $w$ is not unique, since the reduced decomposition of the element $w$ is not unique.
Denote by $l_{C}(w)$ the smallest value $k$ corresponding to any reduced decomposition (1.11). The corresponding set of roots $\{\alpha_{1},\alpha_{2},\dots,\alpha_{k}\}$ consists of linearly independent and not necessarily simple roots, see Lemma 1.2. If $l(w)$ is the smallest value $k$ in (1.11) such that all roots $\alpha_{i}$ are simple, then $l_{C}(w)\leq l(w)$.
Lemma 1.2.
[Ca72, Lemma 3] Let $\alpha_{1},\alpha_{2},\dots,\alpha_{k}\in\varPhi$. Then, $s_{\alpha_{1}}s_{\alpha_{2}}\dots s_{\alpha_{k}}$ is reduced if and only if $\alpha_{1},\alpha_{2},\dots,\alpha_{k}$ are linearly independent. ∎
The admissible diagram may contain cycles, since the roots of $S$ are non necessarily simple, [St17]. Let us fix some basis of roots corresponding to the given admissible diagram $\Gamma$:
$S=\{\alpha_{1},\dots,\alpha_{k},\beta_{1},\dots,\beta_{h}\}$ (1.12)
According to (1.4(a)(a)), there exists a certain set (1.12) of linearly independent roots.Thanks to (1.4(b)(b)), there exists a partition $S=S_{\alpha}\coprod S_{\beta}$ which is said to be the bicolored partition. The admissible diagram is bicolored, i.e., the set of nodes can be partitioned into two disjoint subsets $S_{\alpha}=\{\alpha_{1},\dots,\alpha_{k}\}$ and $S_{\beta}=\{\beta_{1},\dots,\beta_{h}\}$, where roots of $S_{\alpha}$ (resp. $S_{\beta}$) are mutually orthogonal.
Let $w=w_{1}w_{2}$ be the decomposition of $w$ into the product of two involutions. By [Ca72, Lemma 5] each of $w_{1}$ and $w_{2}$ can be expressed as a product of reflections as follows:
$w=w_{1}{w}_{2},\quad\text{ where }\quad w_{1}=s_{\alpha_{1}}s_{\alpha_{2}}\dots s_{\alpha_{k}},\quad w_{2}=s_{\beta_{1}}s_{\beta_{2}}\dots s_{\beta_{h}},$ (1.13)
where subset $S_{\alpha}=\{\alpha_{1},\dots,\alpha_{k}\}$ (resp. $S_{\beta}=\{\beta_{1},\dots,\beta_{h}\}$) consists of mutually orthogonal roots. The decomposition (1.13) is said to be a bicolored decomposition. The subset of roots corresponding to $w_{1}$ (resp. $w_{2}$) is said to be $\alpha$-set (resp. $\beta$-set):
$\alpha\text{-set}=\{\alpha_{1},\alpha_{2},\dots,\alpha_{k}\},\quad\beta\text{-set}=\{\beta_{1},\beta_{2},\dots,\beta_{h}\}.$ (1.14)
The element $w$ in (1.13) is said to be semi-Coxeter element; it represents the conjugacy class associated with the admissible diagram $\Gamma$ and root subset $S$.
Any admissible diagram $\Gamma$ is said to be a Carter diagram if any edge connecting a pair of roots $\{\alpha,\beta\}$ with inner product $(\alpha,\beta)>0$ (resp. $(\alpha,\beta)<0$) is drawn as dotted (resp. solid) edge. There is no edge for the inner product $(\alpha,\beta)=0$.
Up to solid/dotted edges, the classification of Carter diagrams coincides with the classification of admissible diagrams.
1.8. Linkage root and linkage diagram
We consider the extension of the root subset $S$ from the bicolored partition (1.12) by means of the root $\gamma\in\varPhi$, so that the set of roots
$S^{\prime}=\{\alpha_{1},\dots,\alpha_{k},\beta_{1},\dots,\beta_{h},\gamma\}$
is linearly independent. The new diagram $\Gamma^{\prime}$ is obtained by addition new edges, these edges are solid (resp. dotted) for $(\gamma,\tau)=-1$ (resp. $(\gamma,\tau)=1)$, where $\tau\in S$.
The diagram $\Gamma^{\prime}$ is said to be a linkage diagram and the root $\gamma$ is said to be a linkage root. The roots $\tau$ corresponding to the new edges ($(\gamma,\tau)\neq 0$) are said to be endpoints of the linkage diagram. Consider vector $\gamma^{\nabla}$ defined by (1.15). This vector is said to be the linkage label vector. There is a one-to-one correspondence between linkage label vectors $\gamma^{\nabla}$ (with labels $\gamma^{\nabla}_{i}\in\{-1,0,1\}$) and linkage diagrams such that $(\gamma,\tau)\in\{-1,0,1\}$. $\gamma^{\nabla}:=\left(\begin{array}[]{c}(\gamma,\alpha_{1})\\ \dots,\\ (\gamma,\alpha_{k})\\ (\gamma,\beta_{1})\\ \dots,\\ (\gamma,\beta_{h})\\ \end{array}\right)$ (1.15)
1.9. Solid and dotted edges of linkage diagrams
Lemma 1.3.
[St17, Lem. A.1] Every cycle in the Carter diagram contains at least one solid edge and at least one dotted edge.
Figure 2. Example of linkage diagram for $E_{6}(a_{1})$.
This lemma is also valid for cycles of any linkage diagram. In Fig. 2 if $\{\beta_{1},\gamma\}$ is solid then edge $\{\alpha_{3},\gamma\}$ is necessarily dotted, see triangle $\{\alpha_{3},\gamma,\beta_{1}\}$. Similarly, from triangle $\{\alpha_{2},\gamma,\beta_{1}\}$ the edge $\{\alpha_{2},\gamma\}$ is necessarily dotted.
1.10. Getting rid of triangles
By adding the linkage root to a Carter diagram, we get a new diagram which is not necessarily a Carter diagram. After this, we act by the Weyl group of the inverse quadratic form and obtain the orbit of the linkage diagrams. In the most of cases, this orbit contains a Carter diagram, see Figs. 5910. Sometimes the orbit does not contain a Carter diagram because every linkage diagram on the orbit contains a triangle or other odd cycle. By changing the basis we can eliminate a triangle or other odd cycle. The procedure for eliminating the triangle is shown in Fig. 3.
Figure 3. Eliminating the triangle
1.11. Nilpotent orbits in the simple Lie algebras
There is a connection (which was pointed out to me by A. Elashivili) between the representatives of nilpotent orbits in simple Lie algebras and root subsets of linearly independent roots represented by Carter diagrams. By theorem of Jacobson-Morozov, this connection is also a connection between simple three-dimensional $S$-subalgebras (or $sl2$-triples) listed by Dynkin, and root subsets of linearly independent roots, see [Elk72], [Gr08], [D52], [On00], [E75]. [MK21] .
Acknowledgements
I thank A.Elashvili and M.Jibladze for their attention to this paper and useful discussions during its preparation.
2. The main results
2.1. When a vector is a linkage root?
According with §1.8, the linkage diagram is obtained from a Carter diagram $\Gamma$ by adding some linkage root $\gamma$, with its bonds, so that the roots corresponding to vertices of $\Gamma$ together with $\gamma$ form a linearly independent root subset. We would like to describe such roots in the general case, namely:
[…]
With every linkage diagram we associate the linkage label vector. There is one-to-one correspondence between linkage label vectors and linkage diagrams, see §1.8. The linkage diagram and their linkage label vector $\gamma^{\nabla}$ for the Carter diagram $E_{6}(a_{1})$ is depicted in Fig. 2. For more examples of linkage diagrams and linkage label vectors see Fig. 5 (case $D_{4}(a_{1})$), Figs. 10 (case $D_{k}(a_{l})$).
It turns out that the answer to question (2.1) is very simple in terms of the quadratic form associated with the partial Cartan matrix. Let $\gamma$ be a linkage root for $\Gamma$, $\gamma^{\nabla}$ be the linkage label vector, see (1.15). We denote by $L$ the space spanned by roots of $S$, and by $L(\gamma)$ the space spanned by $L$ and $\gamma$. Let us write
$L=[\tau_{1},\dots,\tau_{l}],\qquad L(\gamma)=[\tau_{1},\dots,\tau_{l},\gamma].$ (2.2)
Note that $L(\gamma)$ corresponds to a certain linkage diagram. Let $\mathscr{B}_{\Gamma}$ be the quadratic form associated with the partial Cartan matrix $B_{\Gamma}$. Then,
$\mathscr{B}_{\Gamma}(\gamma)=\langle B_{\Gamma}\gamma,\gamma\rangle=\langle\gamma^{\nabla},B_{\Gamma}^{-1}\gamma^{\nabla}\rangle,$
where $\langle\cdot,\cdot\rangle$ is the usual dot product. Since $B_{\Gamma}$ is positive definite, the eigenvalues of $B_{\Gamma}$ are positive. Therefore, the matrix $B^{-1}_{\Gamma}$ is also positive definite. We call the quadratic form $\mathscr{B}^{\vee}_{\Gamma}$ corresponding to the matrix $B^{-1}_{\Gamma}$ the inverse quadratic form. The form $\mathscr{B}^{\vee}_{\Gamma}$ is positive definite.
2.2. Criterion that a vector is a linkage root
Theorem (Theorem 3.3, Linkage root criterion).
(i) Let $\theta^{\nabla}$ be the linkage label vector corresponding to a certain root $\theta\in L(\gamma)$. The root $\theta$ is a linkage root if and only if
$\mathscr{B}^{\vee}_{\Gamma}(\theta^{\nabla})<2.$ (2.3)
(ii) Let $\theta\in L(\gamma)$ be a root with only one endpoint $\tau_{i}\in S$. Then $\theta$ is a linkage root if and only if
$b^{\vee}_{i,i}<2,$
where $b^{\vee}_{i,i}$ is the $i$th diagonal element of $B^{-1}_{\Gamma}$.
Note that criterion (2.3) does not depend on the choice of the $\Gamma$-set $S$, but only on the quadratic form $\mathscr{B}$ or, what the same, on the diagram $\Gamma$.
Let $\{\Gamma^{\prime},\Gamma\}$ be a pair of Carter diagrams from the list (1.8), $M$ be the transition matrix transforming $\Gamma^{\prime}$-set $S^{\prime}$ to $\Gamma$-set $S$, see §1.6, and let $B_{\Gamma^{\prime}}$ and $B_{\Gamma}$ be the partial Cartan matrices for $\Gamma^{\prime}$ and $\Gamma$. Then, $M$ relates the partial Cartan matrices $B_{\Gamma^{\prime}}$ and $B_{\Gamma}$ as follows:
$^{t}{M}\cdot B_{\Gamma^{\prime}}\cdot M=B_{\Gamma}.\\$
It is shown in Proposition 3.4 that vector $\gamma^{\nabla}$ is the linkage label vector for $\Gamma$ if and only if   $^{t}{M}\gamma^{\nabla}$ is the linkage label vector for $\Gamma^{\prime}$.
2.3. Full linkage systems $\mathscr{L}(D_{l})$ and $\mathscr{L}(D_{l}(a_{k}))$
Denote by $\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime},S^{\prime})$ the set of linkage diagrams
$\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime},S^{\prime})=\{\gamma^{\nabla}\mid\gamma\in\varPhi(\widetilde{\Gamma}),\gamma\not\in L(S^{\prime})\},$ (2.4)
[…]
$\mathscr{L}(\Gamma^{\prime})=\bigcup_{\Gamma^{\prime}\prec\widetilde{\Gamma}}\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime}).$ (2.5)
For instance, for $\Gamma=A_{7}$, we have:
$\begin{split}&\mathscr{L}(A_{7})=\mathscr{L}_{A_{8}}(A_{7})\cup\mathscr{L}_{D_{8}}(A_{7})\cup\mathscr{L}_{E_{8}}(A_{7}).\end{split}$
If the diagram $\widetilde{\Gamma}$ in (2.4) is the $A$-type (resp. $D$-type, resp. $E$-type) Dynkin diagram, the component $\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime})$ is called the $A$-component (resp. $D$-component, resp. $E$-component) in the full linkage system (2.5). In Theorem 5.3, the $D$- and $E$-components, the size of each component for the full linkage systems $\mathscr{L}(D_{l})$ and $\mathscr{L}(D_{l}(a_{k}))$ are found. The results are shown in Table 2.1.
$\Gamma$ Number of Number of linkage diagrams,
components and $p=\mathscr{B}^{\vee}_{\Gamma}(\gamma^{\nabla})$
$D$-components $E$-components   In all
$p=1$ $p=\frac{l}{4}$
$D_{4}$,  $D_{4}(a_{1})$ 33 $3\times 8=24$ - $24$
$D_{5}$,  $D_{5}(a_{1})$ 33 $10$ $2\times 16=32$ $42$
$D_{6}$, $D_{6}(a_{1})$, $D_{6}(a_{2})$ 33 $12$ $2\times 32=64$ $76$
$D_{7}$, $D_{7}(a_{1})$, $D_{7}(a_{2})$ 33 $14$ $2\times 64=128$ $142$
$D_{l}$,  $D_{l}(a_{k})$, $l>7$ 11 $2l$ - $2l$
Table 2.1. Number of linkage diagrams in extensions for diagrams $D_{l}$ and $D_{l}(a_{k})$
For any pair of homogeneous Carter diagrams $\Gamma$ and $\Gamma^{\prime}$ the sizes of linkage systems $\mathscr{L}(\Gamma)$ and $\mathscr{L}(\Gamma^{\prime})$ are the same (Proposition 4.3): $\mid\mathscr{L}(\Gamma)\mid\;\;=\;\;\mid\mathscr{L}(\Gamma^{\prime})\mid.$
The linkage systems $\mathscr{L}(D_{4})$ (resp. $\mathscr{L}(D_{4}(a_{1}))$, resp. $\mathscr{L}(D_{l}$) for $l>4$ , resp. $\mathscr{L}(D_{l}(a_{k}))$ for $l>4$) are depicted in Fig. 4 (resp. Fig. 5, resp. Fig. 9, resp. Fig. 10).
One of two orbits of the $E$-component of the linkage system $\mathscr{L}(D_{5})$ are depicted in Fig. 11. Two $E$-components of the linkage system $\mathscr{L}(D_{6})$ are depicted in Fig. 12.
2.4. The Weyl group of quadratic form $\mathscr{B}^{\vee}_{\Gamma}$
The Weyl group of $\mathscr{B}^{\vee}_{\Gamma}$ is generated by dual reflections $\{s^{*}_{\tau_{i}}\mid\tau_{i}\in S\}$, where
$s^{*}_{\tau_{i}}{\gamma}^{\nabla}:={\gamma}^{\nabla}-\langle{\gamma}^{\nabla},\tau_{i}\rangle\tau_{i}^{\nabla}.$
If $\gamma$$^{\nabla}$ is a linkage label vector then values $(s^{*}_{\tau_{i}}{\gamma}^{\nabla})_{\tau_{k}}$ in (3.20) belong to the set $\{-1,0,1\}$, i.e., we get new linkage label vector $s^{*}_{\tau_{i}}{\gamma}^{\nabla}$ (Proposition 3.7). The action of reflections $s_{\tau_{i}}$ and dual reflections $s^{*}_{\tau_{i}}$ related as follows:
$(s_{\tau_{i}}\gamma)^{\nabla}=(s^{*}_{\tau_{i}}\gamma^{\nabla}).$
The dual reflections $s^{*}_{\tau_{i}}$ preserve the inverse quadratic form $\mathscr{B}^{\vee}_{\Gamma}$ (Proposition 3.9):
$\mathscr{B}^{\vee}_{\Gamma}(s^{*}_{\tau_{i}}{\gamma}^{\nabla})=\mathscr{B}^{\vee}_{\Gamma}({\gamma}^{\nabla})\text{ for }{\gamma}^{\nabla}\in L^{\nabla}.$
The rational number $p=\mathscr{B}^{\vee}_{\Gamma}({\gamma}^{\nabla})$ is the invariant characterizing the given orbit, since all elements of this orbit have the same value $\mathscr{B}^{\vee}_{\Gamma}({\gamma}^{\nabla})$, see Table 2.1.
2.5. Loctets
A spindle-like subset consisting of $8$ linkage label vectors is called a loctet (linkage octet).
The $D$-component of $\mathscr{L}(D_{4})$ consists of three orbits, each orbit consists of one loctet, Fig. 4.
The $D$-component of $\mathscr{L}(D_{5})$ consists of three orbits, two of which consist of one loctet, Fig. 5.
The $E$-component of $\mathscr{L}(D_{5})$ consists of two orbits, each orbit consists of two loctets, Fig. 11.
The $E$-component of $\mathscr{L}(D_{6})$ consists of two orbits, each orbit consists of four loctets, Fig. 12.
The $E$-component of $\mathscr{L}(D_{7})$ consists of two orbits, each orbit consists of eight loctets, the corresponding figures can be found in [St14, Figs. C.63, C.64].
Figure 4. Three components of the linkage system $\mathscr{L}(D_{4})$ contain $24$ elements. The linkage label vectors are shown above, the corresponding linkage diagrams are below.
Figure 5. Three components of the linkage system $\mathscr{L}(D_{4}(a_{1}))$ contain $24$ elements. The linkage label vectors are shown above, the corresponding linkage diagrams are below.
3. Quadratic forms associated with Carter diagrams
Let $L$ be the subspace spanned by the vectors $S=\{\tau_{1},\dots,\tau_{n}\}$. We write this fact as follows:
$L=[\tau_{1},\dots,\tau_{n}],$
The subspace $L$ is said to be the $S$-associated subspace. Let $B$ be the Cartan matrix corresponding to the root system $\varPhi$.
Proposition 3.1.
(i) The restriction of the bilinear form associated with the Cartan matrix $B$ on the subspace $L$ coincides with the bilinear form associated with the partial Cartan matrix $B_{\Gamma}$, i.e., for any pair of vectors $v,u\in L$, we have
$(v,u)_{\stackrel{{\scriptstyle}}{{\Gamma}}}=(v,u),\text{ and }\mathscr{B}_{\Gamma}(v)=\mathscr{B}(v).$ (3.1)
(ii) For every Carter diagram, the matrix $B_{\stackrel{{\scriptstyle}}{{\Gamma}}}$ is positive definite.
Proof. (i) From (1.5) we deduce:
$(v,u)_{\stackrel{{\scriptstyle}}{{\Gamma}}}=(\sum\limits_{i}{t_{i}{\tau_{i}}},\sum\limits_{j}{q_{j}{\tau_{j}}})_{\stackrel{{\scriptstyle}}{{\Gamma}}}=\sum\limits_{i,j}t_{i}{q}_{j}(\tau_{i},\tau_{j})_{\stackrel{{\scriptstyle}}{{\Gamma}}}=\sum\limits_{i,j}t_{i}{q}_{j}(\tau_{i},\tau_{j})=(v,u).$
(ii) This follows from (i). ∎
If $\Gamma$ is a Dynkin diagram, the partial Cartan matrix $B_{\Gamma}$ is the Cartan matrix associated with $\Gamma$. By (3.1) the matrix $B_{\Gamma}$ is positive definite. The symmetric bilinear form associated with $B_{\Gamma}$ is denoted by $(\cdot,\cdot)_{\Gamma}$ and the corresponding quadratic form is denoted by $\mathscr{B}_{\Gamma}$.
3.1. The inverse quadratic form $\mathscr{B}^{\vee}_{\Gamma}$
3.1.1. Action of $B^{-1}_{\Gamma}$ on the space $L$
Let $S=\{\tau_{1},\dots,\tau_{l}\}$ be a $\Gamma$-set, $L$ - the space spanned by roots of $S$, $\gamma$ be a root in $L$:
$\gamma=t_{1}\tau_{1}+\dots+t_{l}\tau_{l},$
where $t_{i}$ are some rational numbers. Then, for $\gamma\in L$, we have
$\gamma^{\nabla}=\left(\begin{array}[]{c}(\gamma,\tau_{1})\\ \dots\\ (\gamma,\tau_{l})\\ \end{array}\right)=\left(\begin{array}[]{c}\sum t_{i}(\tau_{i},\tau_{1})\\ \dots\\ \sum t_{i}(\tau_{i},\tau_{l})\\ \end{array}\right)=B_{\Gamma}\left(\begin{array}[]{c}t_{1}\\ \dots\\ t_{l}\\ \end{array}\right)=B_{\Gamma}\gamma,\text{ and }\gamma={B}^{-1}_{\Gamma}\gamma^{\nabla}.$ (3.2)
3.1.2. The projection of the linkage root
Let $L^{\perp}$ be the orthogonal complement of $L$ to $L(\gamma)$ in the sense of the symmetric bilinear form $(\cdot,\cdot)$ associated with the root system $\varPhi$:
$L(\gamma)=L\oplus L^{\perp}.$
Let $\gamma_{L}$ be the projection of the linkage root $\gamma$ on $L$. For any root $\theta\in L(\gamma)$ such that $\theta\not\in L$, we have $L(\theta)=L(\gamma)$, and $\theta$ is uniquely decomposed into the following sum:
$\theta=\theta_{L}+\mu,\quad\text{ where }\quad\theta_{L}\in L,\quad\mu\in L^{\perp}.$ (3.3)
Given any vector $\theta$ by decomposition (3.3), we introduce also the conjugate vector $\overline{\theta}$ as follows:
$\theta=\theta_{L}+\mu,\quad\overline{\theta}=\theta_{L}-\mu,\quad\text{ where }\quad\theta_{L}\in L,\quad\mu\in L^{\perp}.$
Figure 6.   The roots $\gamma=\gamma_{L}+\mu$ and $\overline{\gamma}=\gamma_{L}-\mu$.
The vector $\mu$ is said to be the normal extending vector, see Fig. 6. The following proposition collects a number of properties of the projection $\gamma_{L}$ of the linkage root $\gamma$, the linkage label vector $\gamma^{\nabla}$, and the normal extending vector $\mu$:
Proposition 3.2.
(i) The linkage label vector $\gamma^{\nabla}$ and the projection $\gamma_{L}$ are related as follows:
$\gamma^{\nabla}={\overline{\gamma}}^{\nabla}=\gamma_{L}^{\nabla}=B_{\Gamma}\gamma_{L}.$ (3.4)
(ii) The vector $\mu$ is, up to sign, a fixed vector for any root $\L\theta\in L(\gamma)\backslash$ L.
(iii) The value $\mathscr{B}_{\Gamma}(\theta_{L})$ is constant for any root $\L\theta\in L(\gamma)\backslash$ L.
(iv) The vector $\theta=\theta_{L}+\mu$ is a root if and only if the conjugate vector $\overline{\theta}=\theta_{L}-\mu$ is a root. In addition, $\L\theta\in L(\gamma)\backslash$ L if and only if $\L\overline{\theta}\in L(\overline{\gamma})\backslash$ L.
(v) Let $\theta$ is a root, $\theta\not\in L$. Then, vector $\theta_{L}+t\mu$ is a root if and only if $t=\pm{1}$.
(vi) If $\delta$ is a root in $\LL(\gamma)\backslash L$ such that $\delta^{\nabla}=\gamma^{\nabla}$, then $\delta=\gamma$ or $\delta=\overline{\gamma}$.
(vii) For any root $\L\theta\in L(\gamma)\backslash$ L, we have
$\mathscr{B}_{\Gamma}^{\vee}(\theta^{\nabla})=\mathscr{B}_{\Gamma}(\theta_{L}),$
and, $\mathscr{B}^{\vee}_{\Gamma}(\theta^{\nabla})$ is a constant for all roots $\L\theta\in L(\gamma)\backslash$ L.
Proof. (i) Let $\gamma=\gamma_{L}+\mu.$ Since $(\mu,\tau_{i})=0$ for any $i$, we have as follows:
$(\gamma,\tau_{i})=(\gamma_{L},\tau_{i})\;\text{ for any }i,\,\text{ i.e., }\gamma^{\nabla}=\gamma_{L}^{\nabla}.$
Then, by (3.2) we have
$\gamma^{\nabla}={\overline{\gamma}}^{\nabla}=\gamma_{L}^{\nabla}=B_{\Gamma}\gamma_{L}.$
(ii) Any root $\theta\in L(\gamma)$ such that $\theta\not\in L$ decomposes as $\tau+\gamma$ for some $\tau\in L$. Since $\gamma=\gamma_{L}+\mu$, where $\gamma_{L}\in L$, we have $\theta=\tau+\gamma_{L}+\mu$. Here, $\tau+\gamma_{L}\in L$. Let us put $\theta_{L}=\tau+\gamma_{L}$. Then $\theta=\theta_{L}+\mu$.
(iii) Let $\mathscr{B}$ be the quadratic form associated with the root system $\varPhi$. By (3.3) we have $\theta_{L}\perp\mu$, and $\mathscr{B}(\theta)=\mathscr{B}(\theta_{L})+\mathscr{B}(\mu)$. Here, $\mathscr{B}(\theta)=2$ since $\theta$ is the root, and by (ii), we have $\mathscr{B}(\theta_{L})=2-\mathscr{B}(\mu)$, i.e, $\mathscr{B}(\theta_{L})$ is constant. By (3.1), we have $\mathscr{B}_{\Gamma}(\theta_{L})=\mathscr{B}(\theta_{L})$, i.e., $\mathscr{B}_{\Gamma}(\theta_{L})$ is also constant for all $\theta\in L(\gamma)$.
(iv) Let $\theta$ be a root, i.e., $\mathscr{B}(\theta)=\mathscr{B}(\theta_{L})+\mathscr{B}(\mu)=2$. Then for $\overline{\theta}$, we have $\mathscr{B}(\overline{\theta})=\mathscr{B}(\theta_{L})+\mathscr{B}(-\mu)=2$ as well. By (1.3) $\overline{\theta}$ is a root as well. Further, $\theta\in L(\gamma)$ means that $\theta=\gamma+\tau$, where $\tau\in L$. Thus, $\theta=\gamma_{L}+\mu+\tau$. Put $\theta_{L}=\gamma_{L}+\tau$. Then $\theta_{L}\in L$, and
$\theta=\theta_{L}+\mu,\;\;\overline{\theta}=\theta_{L}-\mu=\gamma_{L}+\tau-\mu=\overline{\gamma}+\tau\in L(\overline{\gamma}).$
(v) Consider root $\theta=\theta_{L}\pm\mu$. Since $\theta_{L}\perp\mu$, the following relations hold:
$\mathscr{B}(\theta_{L}+t\mu)=\mathscr{B}(\theta\pm\mu+t\mu)=\mathscr{B}(\theta+(t\pm{1})\mu)=\mathscr{B}(\theta)+(t\pm{1})^{2}\mathscr{B}(\mu).$
Because $\theta$ is a root, then $\mathscr{B}(\theta_{L}+t\mu)=\mathscr{B}(\theta)=2$ and $t=\pm{1}$.
(vi) By $\delta^{\nabla}=\gamma^{\nabla}$ and (3.4), the roots $\delta$ and $\gamma$ have the same projections on $L$: $\delta_{L}=\gamma_{L}$. If $\delta=\delta_{L}+t\mu=\gamma_{L}+t\mu$ is a root, then by (v) we have $t=\pm{1}$. Thus, $\delta=\gamma$ or $\overline{\gamma}$.
(vii) By heading (i), since $\theta_{L}\in L$, we have $\theta_{L}^{\nabla}=B_{\Gamma}\theta_{L}$. Thus,
$\mathscr{B}^{\vee}_{\Gamma}(\theta^{\nabla})=\langle B_{\Gamma}^{-1}\theta^{\nabla},\theta^{\nabla}\rangle=\langle{\theta_{L}},B_{\Gamma}\theta_{L}\rangle=\mathscr{B}_{\Gamma}(\theta_{L}).\\$
3.2. The linkage root theorem
In this section we give a criterion for a given vector to be a linkage root.
Theorem 3.3 (Linkage root criterion).
(i) Let $\theta^{\nabla}$ be the linkage label vector corresponding to a certain root $\theta\in L(\gamma)$, i.e., $\theta^{\nabla}=B_{\Gamma}\theta_{L}$. The root $\theta$ is a linkage root, (i.e., $\theta$ is linearly independent of roots of $L$) if and only if
$\mathscr{B}^{\vee}_{\Gamma}(\theta^{\nabla})<2.$ (3.5)
(ii) Let $\theta\in L(\gamma)$ be a root connected only with one $\tau_{i}\in S$. The root $\theta$ is a linkage root if and only if
$b^{\vee}_{ii}<2,$ (3.6)
where $b^{\vee}_{ii}$ is the $i$th diagonal element of $B^{-1}_{\Gamma}$.
Proof. (i) By Proposition 3.2(vi) and Proposition 3.1 we have
$\mathscr{B}^{\vee}_{\Gamma}(\theta^{\nabla})=\mathscr{B}_{\Gamma}(\theta_{L})=\mathscr{B}(\theta_{L})=\mathscr{B}(\theta)-\mathscr{B}(\mu)\leq 2.$ (3.7)
Then, $\mathscr{B}^{\vee}_{\Gamma}(\theta^{\nabla})=2$ if and only if $\mathscr{B}(\mu)=0$. If $\mathscr{B}^{\vee}_{\Gamma}(\theta^{\nabla})=2$ then $\mathscr{B}(\mu)=0$, $\mu=0$ and $\theta=\theta_{L}$. Therefore, $\theta$ is linearly depends on vectors of $L$. If $\mathscr{B}^{\vee}_{\Gamma}(\theta^{\nabla})<2$ then $\mathscr{B}(\mu)>0$, i.e., $\mu\neq 0$ and $\theta$ is linearly independent of roots of $L$.
(ii) We have
$\theta^{\nabla}\quad=\quad\left(\begin{array}[]{c}(\theta,\tau_{1})\\ \dots\\ (\theta,\tau_{i})\\ \dots\\ (\theta,\tau_{l})\\ \end{array}\right)\quad=\quad\left(\begin{array}[]{c}0\\ \dots\\ \pm{1}\\ \dots\\ 0\\ \end{array}\right),$
and $\mathscr{B}^{\vee}_{\Gamma}(\theta^{\nabla})=b^{\vee}_{ii}$. Thus, statement (ii) follows from (i). ∎
3.3. Quadratic forms for homogeneous Carter diagrams
Proposition 3.4.
Let $\{\Gamma^{\prime},\Gamma\}$ be a pair of Carter diagrams from the adjacency list (1.8), $M$ be the transition matrix transforming $\Gamma^{\prime}$-set $S^{\prime}$ to $\Gamma$-set $S$ as in §1.6. The vector $\gamma^{\nabla}$ is the linkage label vector for $\Gamma$ if and only of   $^{t}{M}\gamma^{\nabla}$ is the linkage label vector for $\Gamma^{\prime}$.
Proof. By Theorem 1.1, the transition matrix $M$ maps the $\Gamma^{\prime}$-set $S^{\prime}$ to the $\Gamma$-set $S$:
$M\tau^{\prime}_{i}=\tau_{i},\;\text{ where }\;\tau^{\prime}_{i}\in S^{\prime},\tau_{i}\in S.$
Let $B_{\Gamma^{\prime}}$ and $B_{\Gamma}$ be the partial Cartan matrices for $\Gamma^{\prime}$ and $\Gamma$. The transition matrix $M$ transforms $S^{\prime}$ to $S$ and relates the partial Cartan matrices $B_{\Gamma^{\prime}}$ and $B_{\Gamma}$ as follows:
$^{t}{M}\cdot B_{\Gamma^{\prime}}\cdot M=B_{\Gamma},\\$ (3.8)
see [St23, §4.7]. Since $M$ is an involution, we get
$M\cdot B^{-1}_{\Gamma^{\prime}}\cdot{}^{t}M=B^{-1}_{\Gamma}.$
Further,
$\begin{split}\mathscr{B}^{\vee}_{\Gamma}(\gamma^{\nabla})=&\langle B^{-1}_{\Gamma}\gamma^{\nabla},\gamma^{\nabla}\rangle=\langle B^{-1}_{\Gamma^{\prime}}\cdot{}^{t}M\gamma^{\nabla},{}^{t}M\gamma^{\nabla}\rangle=\mathscr{B}^{\vee}_{\Gamma^{\prime}}({}^{t}M\gamma^{\nabla}).\end{split}$
Then, $\mathscr{B}^{\vee}_{\Gamma}(\gamma^{\nabla})<2$   if and only if   $\mathscr{B}^{\vee}_{\Gamma^{\prime}}({}^{t}M\gamma^{\nabla})<2$. By Theorem 3.3, we get what we need. ∎
3.4. The chain of homogeneous pairs
Consider any homogeneous pair of Carter diagrams $\{\Gamma^{\prime},\Gamma\}$. There exists the chain of homogeneous pairs from the adjacency list (1.8) which connect $\Gamma^{\prime}$ and $\Gamma$ as follows:
$\{\{\Gamma_{0},\Gamma_{1}\},\{\Gamma_{1},\Gamma_{2}\},\dots,\{\Gamma_{k-1},\Gamma_{k}\},\text{ where }\Gamma_{0}=\Gamma^{\prime}\text{ and }\Gamma_{k}=\Gamma.\\$ (3.9)
Let $S_{j}=\{\tau^{j}_{1},\dots,\tau^{j}_{n}\}$ be the associated $\Gamma_{j}$-set, $M^{(j)}$ be the transition matrices connecting $j$th pair in (3.9), where $j=0,\dots k-1$, see Theorem 1.1. Then,
[…]
Two Carter diagrams obtained from each other by a sequence of reflections (1.9), are called similar Carter diagrams, see [St23, Fig. 4]. In fact, mappings $L_{\tau_{i}}$ can be considered as a degenerated case of transition matrices $M$ of Theorem 1.1. The Carter diagrams in the adjacency list (1.8) are given up to similarity, so using the transition matrices $M^{(j)}$ together with mappings $L_{\tau_{i}}$ we can construct the map $F$ transforming the partial Cartan matrices $B_{\Gamma^{\prime}}$ and $B_{\Gamma}$ to each other, also the linkage roots for $\Gamma^{\prime}$ and $\Gamma$ to each other, see [St23, §4.2.1]:
$\begin{split}&F\tau^{\prime}_{i}=\tau_{i},\quad{}^{t}{F}\cdot B_{\Gamma^{\prime}}\cdot F=B_{\Gamma},\quad\mathscr{B}^{\vee}_{\Gamma}(\gamma^{\nabla})=\mathscr{B}^{\vee}_{\Gamma^{\prime}}({}^{t}F\gamma^{\nabla}).\end{split}$ (3.10)
3.5. Spectrum of the Dynkin and Carter diagram
The spectrum of a Dynkin diagram (resp. Carter diagram) is defined as the spectrum of the corresponding Cartan matrix (resp. partial Cartan matrix).
Proposition 3.5.
The spectrum of any Dynkin or Carter diagram lies in the open interval $(0,4)$.
Proof. Note that for any transition matrix $M$ and any similarity matrix $L_{\tau}$ the determinant is equal to 1-1, see (1.10). Then, by (3.8) and (3.10), we have
$\det B_{\Gamma^{\prime}}=\det B_{\Gamma}$ (3.11)
for any homogeneous pair of Carter diagrams $\{\Gamma^{\prime},\Gamma\}$. Let $\Gamma$ be the Dynkin diagram from the corresponding homogeneous class. By (3.11) the spectrum of any Carter diagram coincides with the spectrum of $\Gamma$.
Now consider the spectrum of Dynkin diagrams. By [St08, eq.(3.2)] and [St08, Prop.3.2] eigenvalues $\lambda$ of the Coxeter transformation and the eigenvalues $\rho$ of $B_{\Gamma}$ are related as follows:
$\begin{split}&\frac{(\lambda+1)^{2}}{4\lambda}=(\frac{\rho}{2}-1)^{2}\quad\text{ or }\\ &\lambda+2+\frac{1}{\lambda}=(\rho-2)^{2}.\end{split}$ (3.12)
Since eigenvalues of the Coxeter transformation are roots of unity, see [Bo02], any $\lambda$ is equal to $\cos\varphi+i\sin\varphi$ for some angle $\varphi$. By (3.12)
[…]
Therefore, $0\leq\rho\leq 4$. By Proposition 3.1(ii) $B_{\Gamma}$ is positive definite, then $\rho>0$. For the finite Weyl groups, $1$ is not an eigenvalue of the Coxeter transformation, then $\rho\neq 4$. Thus, $0<\rho<4$. ∎
Using the spectrum of the Cartan matrix could be an invariant description of the Dynkin and Carter diagrams, see [St23, §1.5.3], [MS07].
3.6. The Weyl group of the quadratic form
3.6.1. Dual reflections
Let $\Gamma$ be a Carter diagram, $S=\{\tau_{1},\dots,\tau_{l}\}$ be some $\Gamma$-set, and $L$ be $S$-associated subspace:
$L=[\tau_{1},\dots,\tau_{l}].$
For any $\tau_{i}\in S$, we define vectors $\tau_{i}^{\nabla}$ as follows:
$\tau_{i}^{\nabla}:=B_{\Gamma}{\tau_{i}}.$ (3.13)
Let ${L}^{\nabla}$ be the linear space spanned by the vectors $\tau_{i}^{\nabla}$, where $\tau_{i}\in S$. The map $\tau_{i}\longrightarrow\tau_{i}^{\nabla}$ given by (3.13) is expanded to the linear mapping $L\longrightarrow{L}^{\nabla}$, and
$u^{\nabla}=B_{\Gamma}{u}=\left(\begin{array}[]{c}(u,\tau_{1})\\ \dots\\ (u,\tau_{l})\\ \end{array}\right)\text{ for any }u\in L,$ (3.14)
see (3.2). Consider the restriction of the reflection $s_{\tau_{i}}$ on the subspace $L$. For any $v\in L$, by Proposition 3.1 we have:
$s_{\tau_{i}}v=v-2\frac{(\tau_{i},v)}{(\tau_{i},\tau_{i})}\tau_{i}=v-(\tau_{i},v)_{\stackrel{{\scriptstyle}}{{\Gamma}}}\tau_{i}=v-\langle B_{\Gamma}\tau_{i},v\rangle\tau_{i}=v-\langle\tau_{i}^{\nabla},v\rangle\tau_{i}.$ (3.15)
We define the dual reflection $s^{*}_{\tau_{i}}$ acting on a vector $u^{\nabla}\in L^{\nabla}$ as follows:
$s^{*}_{\tau_{i}}u^{\nabla}:=u^{\nabla}-\langle u^{\nabla},\tau_{i}\rangle\tau_{i}^{\nabla}.$ (3.16)
Let $W_{S}$ (resp. $W_{S}^{\vee}$) be the group generated by reflections $\{s_{\tau_{i}}\mid\tau_{i}\in S\}$ (resp. $\{s^{*}_{\tau_{i}}\mid\tau_{i}\in S\})$, where $S=\{\tau_{1},\dots,\tau_{l}\}$. The group $W_{S}$ is uniquely determined by the quadratic form $\mathscr{B}_{\Gamma}$. In Proposition 3.9 it will be shown that $W_{S}$ (resp. $W_{S}^{\vee}$) preserves $\mathscr{B}_{\Gamma}$ (resp. $\mathscr{B}^{\vee}_{\Gamma}$), so $W_{S}$ (resp. $W_{S}^{\vee}$) is called the Weyl group of $\mathscr{B}_{\Gamma}$ (resp. $\mathscr{B}^{\vee}_{\Gamma}$)111 The Weyl group of the integral quadratic form and their roots were studied by A. Roiter in [R78], and then this study was continued in [GR97], [S18], [MM19], [MZ22] and others..
Proposition 3.6.
(i) For any $\tau_{i}\in S$, we have
$s^{*}_{\tau_{i}}={}^{t}{s}_{\tau_{i}}={}^{t}{s}_{\tau_{i}}^{-1}.$ (3.17)
(ii) The mapping
$\pi:w\rightarrow{}^{t}{w}^{-1}$
determines an isomorphism between $W_{S}$ and $W_{S}^{\vee}$.
Proof. (i) By (3.15) and (3.16), for any $v\in L,u^{\nabla}\in L^{\nabla}$, we have:
$\begin{split}&\langle s^{*}_{\tau_{i}}u^{\nabla},v\rangle=\langle u^{\nabla}-\langle u^{\nabla},\tau_{i}\rangle\tau_{i}^{\nabla},v\rangle=\langle u^{\nabla},v\rangle-\langle u^{\nabla},\tau_{i}\rangle\langle v,\tau_{i}^{\nabla}\rangle,\\ &\langle u^{\nabla},s_{\tau_{i}}v\rangle=\langle u^{\nabla},v-\langle\tau_{i}^{\nabla},v\rangle\tau_{i}\rangle=\langle u^{\nabla},v\rangle-\langle\tau_{i}^{\nabla},v\rangle\langle u^{\nabla},\tau_{i}\rangle.\end{split}$
Thus,
$\langle s^{*}_{\tau_{i}}u^{\nabla},v\rangle=\langle u^{\nabla},s_{\tau_{i}}v\rangle,\quad\text{ for any }v\in L,u^{\nabla}\in L^{\nabla},$
and (3.17) holds.
(ii) Let $\pi(w_{i})={}^{t}w^{-1}_{i}$ for $i=1,2$. Then,
$\pi(w_{1})\,\pi(w_{2})={}^{t}w^{-1}_{1}\;{}^{t}w^{-1}_{2}={}^{t}(w^{-1}_{2}\;w^{-1}_{1})={}^{t}(w_{1}\;w_{2})^{-1}=\pi(w_{1}\;w_{2}).$
Let us put
$w^{*}:={}^{t}{w}^{-1}.$
Then, $\pi(w)=w^{*}$, and
$w^{*}_{1}\;w^{*}_{2}=(w_{1}w_{2})^{*}.$ (3.18)
3.6.2. Action of the Weyl group $W_{S}^{\vee}$
By (3.13), (3.14) and (3.16), for any $u^{\nabla}\in L^{\nabla}$, we have
$\begin{split}&s^{*}_{\tau_{i}}u^{\nabla}=u^{\nabla}-\langle u^{\nabla},\tau_{i}\rangle\tau^{\nabla}_{i}=u^{\nabla}-u^{\nabla}_{\tau_{i}}(\tau^{\nabla}_{i})=u^{\nabla}-u^{\nabla}_{\tau_{i}}(B_{\Gamma}\tau_{i}),\\ &(s^{*}_{\tau_{i}}u^{\nabla})_{\tau_{k}}=u^{\nabla}_{\tau_{k}}-u^{\nabla}_{\tau_{i}}(B_{\Gamma}\tau_{i})_{\tau_{k}}=u^{\nabla}_{\tau_{k}}-u^{\nabla}_{\tau_{i}}(\tau_{i},\tau_{k}).\end{split}$ (3.19)
Let $u$ be a linkage root, then vector $u^{\nabla}$ is the linkage label vector, the coordinates of $u^{\nabla}$ belong to the set $\{-1,0,1\}$. By (3.19) we get
[…]
Proposition 3.7 (Action of dual reflection).
If $u^{\nabla}$ is a linkage label vector then values $(s^{*}_{\tau_{i}}u^{\nabla})_{\tau_{k}}$ in (3.20) belong to the set $\{-1,0,1\}$, i.e., we obtain a new linkage label vector $s^{*}_{\tau_{i}}u^{\nabla}$.
Proof. Let $\{\tau_{k},\tau_{i}\}$ be a solid edge. If $u^{\nabla}_{\tau_{k}}=u^{\nabla}_{\tau_{i}}=1$ (resp. $u^{\nabla}_{\tau_{k}}=u^{\nabla}_{\tau_{i}}=-1$) then roots $\{-u,\tau_{k},\tau_{i}\}$ (resp. $\{u,\tau_{k},\tau_{i}\}$) constitute the root system corresponding to the extended Dynkin diagram $\widetilde{A}_{2}$, which is impossible, see [St17, Lemma A.1]. For remaining pairs $\{u^{\nabla}_{\tau_{k}},u^{\nabla}_{\tau_{i}}\}$, we have $1\leq u^{\nabla}_{\tau_{k}}+u^{\nabla}_{\tau_{i}}\leq 1$. Now, let $\{\tau_{k},\tau_{i}\}$ be a dotted edge. If $u^{\nabla}_{\tau_{k}}=1$ and $u^{\nabla}_{\tau_{i}}=-1$ (resp. $u^{\nabla}_{\tau_{k}}=-1$ and $u^{\nabla}_{\tau_{i}}=1$) then roots $\{u,-\tau_{k},\tau_{i}\}$ (resp. $\{u,\tau_{k},-\tau_{i}\}$) constitute the root system $\widetilde{A}_{2}$, which is impossible. For remaining pairs $\{u^{\nabla}_{\tau_{k}},u^{\nabla}_{\tau_{i}}\}$, we have $1\leq u^{\nabla}_{\tau_{k}}-u^{\nabla}_{\tau_{i}}\leq 1$. ∎
Remark 3.8.
Eq. (3.20) is very useful for calculating linkage systems. Note that action of dual reflections in (3.20) $s^{*}_{\tau_{k}}$ on the linkage label vector $u^{\nabla}$ is non-trivial if and only if $u^{\nabla}_{k}\neq 0$.
Proposition 3.9.
(i) For dual reflections $s^{*}_{\tau_{i}}$, the following relations hold:
$B_{\Gamma}s_{\tau_{i}}=s^{*}_{\tau_{i}}B_{\Gamma},$ (3.21)
$(s_{\tau_{i}}\gamma)^{\nabla}=s^{*}_{\tau_{i}}B_{\Gamma}\gamma_{L}=s^{*}_{\tau_{i}}\gamma^{\nabla}.$ (3.22)
(ii) For $w^{*}\in W^{\vee}_{S}$, we have
$(w\gamma)^{\nabla}=w^{*}{\gamma}^{\nabla}.$
(iii) The following relations hold:
$\begin{split}&\mathscr{B}_{\Gamma}(s_{\tau_{i}}v)=\mathscr{B}_{\Gamma}(v)\text{ for any }v\in L,\\ &\mathscr{B}^{\vee}_{\Gamma}(s^{*}_{\tau_{i}}u^{\nabla})=\mathscr{B}^{\vee}_{\Gamma}(u^{\nabla})\text{ for any }u^{\nabla}\in L^{\nabla}.\end{split}$ (3.23)
Proof. (i) The equality (3.21) is true since for any $u,v\in L$ the following is true:
$\begin{split}&(s_{\tau_{i}}{u},v)_{\Gamma}=({u},s_{\tau_{i}}{v})_{\Gamma},\text{ i.e., }\langle B_{\Gamma}s_{\tau_{i}}{u},v\rangle=\langle B_{\Gamma}{u},s_{\tau_{i}}{v}\rangle=\langle s^{*}_{\tau_{i}}{B}_{\Gamma}{u},v\rangle\text{, and }\\ &\langle(B_{\Gamma}s_{\tau_{i}}-s^{*}_{\tau_{i}}{B}_{\Gamma}){u},v\rangle=0.\end{split}$
Let us consider eq. (3.22). Since $(\tau_{i},\mu)=0$ for any $\tau_{i}\in S$, and $s_{\tau_{i}}\mu=\mu$, we have
[…]
Then, by (3.21) and (3.4)
$(s_{\tau_{i}}\gamma)^{\nabla}=(s_{\tau_{i}}\gamma_{L})^{\nabla}=B_{\Gamma}{s}_{\tau_{i}}\gamma_{L}=s^{*}_{\tau_{i}}B_{\Gamma}\gamma_{L}=s^{*}_{\tau_{i}}\gamma^{\nabla}.$
(ii) Let $w=s_{\tau_{1}}s_{\tau_{2}}\dots s_{\tau_{m}}$ be a decomposition of $w\in W$. Since $s^{*}_{\tau}={}^{t}s^{-1}_{\tau}={}^{t}s_{\tau}$, we deduce from (3.18) and (3.22) the following:
$\begin{split}&(w\gamma)^{\nabla}=(s_{\tau_{1}}s_{\tau_{2}}\dots s_{\tau_{m}}\gamma)^{\nabla}=s^{*}_{\tau_{1}}(s_{\tau_{2}}\dots s_{\tau_{m}}\gamma)^{\nabla}=s^{*}_{\tau_{1}}s^{*}_{\tau_{2}}({s}_{\tau_{3}}\dots s_{\tau_{m}}\gamma)^{\nabla}=\dots=\\ &s^{*}_{\tau_{1}}{s}^{*}_{\tau_{2}}\dots s^{*}_{\tau_{m}}{\gamma}^{\nabla}={w}^{*}{\gamma}^{\nabla}.\end{split}$
(iii) Further, by (3.21) we have
[…]
so (3.23) holds. ∎
By Propositions 3.7 each linkage label vector belongs to some orbit of the group $W^{\vee}_{S}$ action. By Proposition 3.9(iii) all elements of some orbit have the same value $\mathscr{B}^{\vee}_{\Gamma}(u^{\nabla})$. The rational number $p=\mathscr{B}^{\vee}_{\Gamma}(u^{\nabla})$ is the invariant characterizing the given orbit.
To find the preimage root $\gamma_{1}$ for some label vector $\gamma_{1}^{\nabla}$ is suffices to know the preimage for one element $\gamma$ on the orbit. By Proposition 3.9(ii) if $\gamma_{1}^{\nabla}=w^{*}\gamma^{\nabla}$ for some $w$, then $\gamma_{1}=w\gamma$.
4. Extensions
4.1. The vertex extension
Let $\Gamma$ be one of simply-laced Dynkin diagrams, and let $\Gamma^{\prime}$ be one of the Carter diagram out of the homogeneous class $C(\Gamma)$. Each Carter diagram $\Gamma^{\prime}$ from $C(\Gamma)$ has the same rank as ${\rm rank}(\Gamma)$. Consider any $\Gamma$-set $S=\{\tau_{1},\dots,\tau_{n}\}$ and any $\Gamma^{\prime}$-set, $S^{\prime}=\{\tau^{\prime}_{1},\dots,\tau^{\prime}_{n}\}$ obtained by the transformation constructed according to Theorem 1.1. Then the linear spans for $S$ and $S^{\prime}$ coincide:
$L(S):={\rm Span}(S),\quad L(S^{\prime}):={\rm Span}(S^{\prime}),\quad L(S)=L(S^{\prime}).$ (4.1)
[…]
The choose of the root system $\widetilde{\Gamma}$ is ambiguous. The pair $\{\Gamma^{\prime},\widetilde{\Gamma}\}$, where $\Gamma^{\prime}\in C(\Gamma)$, is said to be the vertex extension of $\Gamma^{\prime}$ and is denoted by $\Gamma^{\prime}\prec\widetilde{\Gamma}$.
Figure 7. Two vertex extensions of $D_{5}$: $D_{5}\prec E_{6}$ and $D_{5}\prec D_{6}$.
Remark 4.1.
It may happen that a given root subset $L$ can be extended to two different subsets $L(\gamma)\neq L(\delta)$ such that $\mathscr{B}^{\vee}_{\Gamma}(\gamma^{\nabla})\neq\mathscr{B}^{\vee}_{\Gamma}(\delta^{\nabla})$. For example, there are two extensions for $D_{5}$ with different values of $\mathscr{B}^{\vee}_{\Gamma}(\theta^{\nabla})$:
$\mathscr{B}^{\vee}_{\Gamma}(\gamma^{\nabla})=\frac{5}{4}\quad\text{ for }\quad D_{5}\prec E_{6},\qquad\mathscr{B}^{\vee}_{\Gamma}(\delta^{\nabla})=1\quad\text{ for }\quad D_{5}\prec D_{6},$
see Fig. 7. The matrices $B_{\Gamma}$ and $B_{\Gamma}^{-1}$ for $\Gamma=D_{5}$ are as follows:
[…]
4.2. Partial linkage systems
Let $\widetilde{\Gamma}$ be a Dynkin diagram such that the pair $\{\widetilde{\Gamma},\Gamma\}$ is the vertex extension $\Gamma^{\prime}\prec\widetilde{\Gamma}$, see §4.1. Denote by $\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime},S^{\prime})$ the set of linkage diagrams
$\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime},S^{\prime})=\{\gamma^{\nabla}\mid\gamma\in\varPhi(\widetilde{\Gamma}),\gamma\not\in L(S^{\prime})\},$ (4.2)
[…]
Proposition 4.2.
The set $\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime},S^{\prime})$ does not depend on choosing $\Gamma^{\prime}$-set $S^{\prime}$.
[…]
$\gamma_{1}=\mu_{1}+\sum\limits_{j=1}^{l}t_{j}\tau^{1}_{j},\\$ (4.3)
[…]
$\gamma_{2}=\mu_{2}+\sum\limits_{j=1}^{l}t_{j}\tau^{2}_{j},\\$ (4.4)
where that basis vectors $\tau_{j}^{1}$ and $\tau_{j}^{2}$ have the same coefficient $t_{j}$. Here, $\mu_{1},\mu_{2}$ are normal extending vectors from §3.1.2. Coordinates of the linkage label vectors $\gamma^{\nabla}_{1}$ and $\gamma^{\nabla}_{2}$ are as follows;
$(\gamma^{\nabla}_{1})_{k}=\sum\limits_{j=1}^{l}t_{j}(\tau^{1}_{j},\tau^{1}_{k}),\quad(\gamma^{\nabla}_{2})_{k}=\sum\limits_{j=1}^{l}t_{j}(\tau^{2}_{j},\tau^{2}_{k}),\quad k=1,\dots,l.$ (4.5)
Since inner products $(\tau^{1}_{j},\tau^{1}_{k})$ and $(\tau^{1}_{j},\tau^{1}_{k})$ correspond to the same edge of $\Gamma^{\prime}$ (or both are equal to 0) we get $\gamma^{\nabla}_{1}=\gamma^{\nabla}_{2}$. Thus, we get
$\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime},S^{\prime}_{1})=\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime},S^{\prime}_{2}).$ (4.6)
Thus, the set of linkage diagrams (4.2) can be denoted by $\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime}$). The set of linkage diagrams $\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime})$ is said to be linkage system $\Gamma^{\prime}$ over $\Gamma$ or partial linkage system.
4.3. Full linkage systems
The union of all partial systems by all possible vertex extensions is said to be the linkage system or full linkage system of $\Gamma$:
$\mathscr{L}(\Gamma^{\prime})=\bigcup_{\Gamma^{\prime}\prec\widetilde{\Gamma}}\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime}).$
Proposition 4.3.
(i) For the homogeneous Carter diagrams $\Gamma$ and $\Gamma^{\prime}$ the sizes of the full linkage systems are the same:
$\mid\mathscr{L}(\Gamma)\mid\;\;=\;\;\mid\mathscr{L}(\Gamma^{\prime})\mid.$ (4.7)
(ii) For the full linkage systems (4.7), the estimate of the number of linkage diagrams is as follows:
$\mid\mathscr{L}(\Gamma^{\prime})\mid\;=\;\mid\mathscr{L}(\Gamma)\mid\;\leq\;\bigcup\big{(}\mid\varPhi(\widetilde{\Gamma}_{i})\mid-\mid\varPhi(\Gamma)\mid\big{)},$
where the union is taken by all vertex extensions $\Gamma^{\prime}\prec\widetilde{\Gamma}_{i}$.
Proof. (i) This fact follows from Proposition 3.4 and §3.4.
(ii) By (4.7), for any $\Gamma^{\prime}$ out of the homogeneous class $C(\Gamma)$ we have
$\mid\mathscr{L}_{\widetilde{\Gamma}}(\Gamma^{\prime})\;=\;\mid\mathscr{L}_{\widetilde{\Gamma}}(\Gamma)\mid\;\;\leq\;\;\mid\varPhi(\widetilde{\Gamma})\mid-\mid\varPhi(\Gamma)\mid.$ (4.8)
Then, the number of linkage diagrams in the full linkage systems can be estimated as follows:
$\mid\mathscr{L}(\Gamma^{\prime})\mid\;=\;\mid\mathscr{L}(\Gamma)\mid\;\;\leq\;\;\bigcup\big{(}\mid\varPhi(\widetilde{\Gamma}_{i})\mid-\mid\varPhi(\Gamma)\mid\big{)},$
where the union is taken by all vertex extensions $\Gamma^{\prime}\prec\widetilde{\Gamma}_{i}$. ∎
5. The linkage systems $\mathscr{L}(D_{l})$ and $\mathscr{L}(D_{l}(a_{k}))$.
In this section we assume that $\Gamma$ is one of the Carter diagrams $D_{l}$ or $D_{l}(a_{k})$.
5.1. $D$-components $\mathscr{L}_{D_{l+1}}(D_{l})$ and $\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))$
Lemma 5.1.
Let $D_{l}\prec D_{l+1}$ be some vertex extension with root systems
$\varPhi(D_{l})=\{\tau_{1},\dots,\tau_{l}\},\quad\varPhi(D_{l+1})=\{\tau_{1},\dots,\tau_{l},\tau_{l+1}\},$
where $\tau_{l+1}$ is a simple positive root in $\varPhi(D_{l+1})\backslash\varPhi(D_{l})$. Let $\varphi$ be some positive root in $\varPhi(D_{l+1})\backslash\varPhi(D_{l})$, and $\mu_{max}$ be maximal root in $\varPhi(D_{l+1})$.
(i) The vector
$\delta=\mu_{max}-\varphi+\tau_{l+1}$ (5.1)
is also a root in $\varPhi(D_{l+1})\backslash\varPhi(D_{l})$.
(ii) The linkage label vectors $\varphi^{\nabla}$ and $\delta^{\nabla}$ coincide.
(iii) For any $l\geq 4$, the linkage systems $\mathscr{L}_{D_{l+1}}(D_{l})$ and $\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))$, where $1\leq k\leq\big{[}\frac{l-2}{2}\big{]}$, have $2l$ elements:
$\mid\mathscr{L}_{D_{l+1}}(D_{l})\mid\;=\;\mid\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))\mid\;=\;2l.$ (5.2)
Figure 8.   The vertex extension $D_{l}\prec D_{l+1}$.
Proof. Let $\mathscr{B}$ be the quadratic Tits form associated with $D_{l+1}$.
(i) By (1.3) it suffices to prove that $\mathscr{B}(\delta)=2$. If $\varphi=\mu_{max}$ (resp. $\tau_{l+1}$), then $\delta=\tau_{l+1}$ (resp. $\mu_{max}$). In both cases, $\delta$ is the root in $\varPhi(D_{l+1})\backslash\varPhi(S)$. Suppose $\varphi\neq\mu_{max},\tau_{l+1}$. We need to prove that $\mathscr{B}(\mu_{max}-\varphi+\tau_{l+1})=2$, i.e.,
$\mathscr{B}(\mu_{max})+\mathscr{B}(\varphi)+\mathscr{B}(\tau_{l+1})+2(\mu_{max},\tau_{l+1})-2(\varphi,\mu_{max}+\tau_{l+1})=2.$ (5.3)
Since $\mathscr{B}(\varphi)=\mathscr{B}(\tau_{l+1})=\mathscr{B}(\mu_{max})=2$, eq. (5.3) is equivalent to the following:
$(\varphi,\mu_{max}+\tau_{l+1})-(\mu_{max},\tau_{l+1})=2.$
Further, $(\mu_{max},\tau_{i})=0$ for any $i\neq l$. In particular, $(\mu_{max},\tau_{l+1})=0$ and it suffices to prove that
$(\varphi,\mu_{max}+\tau_{l+1})=2\text{ for any }\varphi\in\varPhi(D_{l+1})\backslash\varPhi(D_{l}).$ (5.4)
We have
[…]
Since $\varphi$ is the positive root in $\varPhi(D_{l+1})\backslash\varPhi(D_{l})$, then $\tau_{l+1}$ appears with a unit coefficient in the decomposition of $\varphi$ in the basis $\{\tau_{1},\dots,\tau_{l+1}\}$. Then, by (5.5) the relation (5.4) holds for any root $\varphi\in\varPhi(D_{l+1})\backslash\varPhi(D_{l})$. Therefore, $\mathscr{B}(\delta)=2$, and $\delta$ is also a root.
(ii) By (5.1) we have
$(\delta,\tau_{i})=(\mu_{max}+\tau_{l+1},\tau_{i})-(\varphi,\tau_{i}),\text{ where }1\leq i\leq l.$
By (5.5) for $i\leq l$, we have
$\delta^{\nabla}_{i}=(\delta,\tau_{i})=-(\varphi,\tau_{i})=-\varphi^{\nabla}_{i},\text{ where }1\leq i\leq l.$
This means that $\delta^{\nabla}=-\varphi^{\nabla}$.
(iii) The number of roots of $\varPhi(D_{l})$ is 2l(l1)2l(l-1), see [Bo02, Table IV]. Then by (4.7) and (4.8) we have
$\mid\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))\mid~=~\mid\mathscr{L}_{D_{l+1}}(D_{l})\mid~\leq~2(l+1)l-2l(l-1)=4l.\\$
By (ii), this size is two times smaller:
$\mid\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))\mid~=~\mid\mathscr{L}_{D_{l+1}}(D_{l})\mid~\leq~2l.$
In Fig. 9, we present exactly $2l$ linkage label vectors for the Carter diagram $D_{l}$. To construct the partial linkage system $\mathscr{L}_{D_{l+1}}(D_{l})$ (resp. $\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))$, we use the action of the Weyl group $W^{\vee}_{S}$, see §3.6.2, and Remark 3.8, see Fig. 10.
So, we conclude that there are exactly $2l$ linkage label vectors in the $D$-component $\mathscr{L}_{D_{l+1}}(D_{l})$. Then, by (4.7) this is also true for the $D$-component $\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))$. ∎
5.2. $E$-component $\mathscr{L}_{E_{8}}(D_{7})$.
By (5.2) the linkage system $\mathscr{L}_{D_{8}}(D_{7})$ contains $14$ elements. We will show that all of them lie in $\mathscr{L}_{E_{8}}(D_{7})$.
Lemma 5.2.
There are $14$ pairs of roots $\eta,\lambda\in\varPhi(E_{8})$ such that $\eta^{\nabla}=-\lambda^{\nabla}$.
Proof. The $7$ pairs of roots $\eta,\lambda\in\varPhi(E_{8})$ such that $\eta^{\nabla}=-\lambda^{\nabla}$ listed in Table 5.2. The opposite roots form $7$ pairs of negative roots $\eta,\lambda\in\varPhi(E_{8})$ such that $\eta^{\nabla}=-\lambda^{\nabla}$. For $1\leq i\leq 7$, the sum $\eta_{i}+\lambda_{i}$ is the same vector, the order of coordinates is shown in parentheses:
$\eta_{i}+\lambda_{i}\;=\;\begin{array}[]{ccccccc}4&7&10&8&6&4&2\\ &&5\\ \end{array},\quad\biggl{(}\begin{array}[]{ccccccc}\tau_{1}&\tau_{3}&\tau_{4}&\tau_{5}&\tau_{6}&\tau_{7}&\tau_{8}\\ &&\tau_{2}\\ \end{array}\biggr{)}$ (5.6)
Vectors $\eta_{i}+\lambda_{i}$ from (5.6) are orthogonal to any $\tau_{i},\;2\leq i\leq 8$, so $\eta_{i}^{\nabla}=-\lambda_{i}^{\nabla}$. Note that there is no need orthogonality for $\tau_{1}$, see (1.15). Further, we calculate the label vectors $\eta_{i}^{\nabla}$ (or $\lambda_{i}^{\nabla}$). They are listed in the last column of Table 5.2.
$\eta\in\varPhi(E_{8})\backslash\varPhi(D_{7})$ $\lambda\in\varPhi(E_{8})\backslash\varPhi(D_{7})$ $\eta^{\nabla}=-\lambda^{\nabla}\in\mathscr{L}_{E_{8}}(D_{7})$
11 $\begin{array}[]{ccccccc}2&3&4&3&2&1&0\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&5&4&3&2\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&0&0&0&0&-1\\ &0\end{array}$
22 $\begin{array}[]{ccccccc}2&3&4&3&2&1&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&5&4&3&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&0&0&0&-1&1\\ &0\end{array}$
33 $\begin{array}[]{ccccccc}2&3&4&3&2&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&5&4&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&0&0&-1&1&0\\ &0\end{array}$
44 $\begin{array}[]{ccccccc}2&3&4&3&3&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&5&3&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&0&-1&1&0&0\\ &0\end{array}$
55 $\begin{array}[]{ccccccc}2&3&4&4&3&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&6&4&3&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}0&-1&1&0&0&0\\ &0\end{array}$
66 $\begin{array}[]{ccccccc}2&3&5&4&3&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}2&4&5&4&3&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}-1&1&0&0&0&0\\ &-1\end{array}$
77 $\begin{array}[]{ccccccc}2&3&5&4&3&2&1\\ &&3\end{array}$ $\begin{array}[]{ccccccc}2&4&5&4&3&2&1\\ &&2\end{array}$ $\begin{array}[]{ccccccc}-1&0&0&0&0&0\\ &1\end{array}$
Table 5.2. The $7$ pairs of positive roots $\eta,\lambda\in\varPhi(E_{8})$ such that $\eta^{\nabla}=-\lambda^{\nabla}$.
5.3. The size and structure of linkage systems $\mathscr{L}(D_{l})$ and $\mathscr{L}(D_{l}(a_{k}))$
Theorem 5.3.
(i) For $l=4$, the full linkage system $\mathscr{L}(D_{4})$ (resp. $\mathscr{L}(D_{4}(a_{1}))$ consists only of $D$-component, see Figs. 45. Each of these $D$-components consists of three orbits of size $8$ each.
(ii) For $l>7$, the full linkage system $\mathscr{L}(D_{l})$ (resp. $\mathscr{L}(D_{l}(a_{1}))$) consists only of $D$-component for any $k$, see Figs. 910. Each of these $D$-components consists of one orbit of size $2l$.
(iii) For $l=5$, the full linkage system $\mathscr{L}(D_{5})$ (resp. $\mathscr{L}(D_{5}(a_{1}))$) consists of the $D$-component ($10$ elements) and the $E$-component consisting of two orbits ($2\times 16$ elements). In total, $\mathscr{L}(D_{5})$ contains $42$ elements. One of two orbits of the $E$-component of the $\mathscr{L}(D_{5})$ is shown in Fig. 11.
(iv) For $l=6$, the full linkage system $\mathscr{L}(D_{6})$ (resp. $\mathscr{L}(D_{6}(a_{1}))$, resp. $\mathscr{L}(D_{6}(a_{2}))$ consists of the $D$-component ($12$ elements) and the $E$-component consisting of two orbits ($2\times 32$ elements). In total, $\mathscr{L}(D_{6})$ contains $76$ elements. Two orbits of the $E$-component of the linkage system $\mathscr{L}(D_{6})$ are shown in Fig. 12.
(v) For $l=7$, the full linkage system $\mathscr{L}(D_{7})$ (resp. $\mathscr{L}(D_{7}(a_{1}))$, resp. $\mathscr{L}(D_{7}(a_{2}))$) consists of one $D$-component ($14$ elements) and two $E$-components consisting of two orbits ($2\times 64$ elements). In total, $\mathscr{L}(D_{7})$ contains $142$ elements.
Proof. (i) Three orbits of $\mathscr{L}_{D_{5}}(D_{4})$ are shown in Fig. 4. Three orbits $\mathscr{L}_{D_{5}}(D_{4}(a_{1}))$ are shown in Fig. 5.
(ii) For $l>7$, the $D$-component $\mathscr{L}_{D_{l+1}}(D_{l})$ and $\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))$ are depicted in Figs. 9 and 10. By Lemma 5.1(iii) there are $2l$ elements in each of these components.
(iii) By (4.7) and (4.8) for $l=5$,
$\mid\mathscr{L}_{E_{6}}(D_{5}(a_{1}))\mid~=~\mid\mathscr{L}_{E_{6}}(D_{6})\mid~\leq~\mid\varPhi(E_{6})\mid-\mid\varPhi(D_{5})\mid~=~72-40=32.\\$
The full linkage system $\mathscr{L}(D_{5})$ consists of $10$-element linkage system $\mathscr{L}_{D_{6}}(D_{5})$ (Lemma 5.1(iii)) and two orbits of the $E$-component, each of which contains $16$ linkages, i.e.,
$\mid\mathscr{L}(D_{5})\mid~=~\mid\mathscr{L}(D_{5}(a_{1}))\mid~\leq~32+10=42.\\$ (5.7)
One of these components is depicted in Fig. 11. Another $E$-component is constructed from the first one as follows: for any linkage diagram the solid edges become dotted, and vice versa, the corresponding label vector becomes just opposite. Thus, in (5.7) both $\mathscr{L}(D_{5})$ and $\mathscr{L}(D_{5}(a_{1}))$ contain exactly $42$ elements.
(iv) Similarly, for $l=6$,
$\mid\mathscr{L}_{E_{7}}(D_{6}(a_{2}))\mid~=~\mid\mathscr{L}_{E_{7}}(D_{6}(a_{1}))\mid~=~\mid\mathscr{L}_{E_{7}}(D_{6})\mid~\leq~\mid\varPhi(E_{7})\mid-\mid\varPhi(D_{6})\mid~\leq~126-60=66.\\$
For the full linkage system, we have the following estimate:
$\mid\mathscr{L}(D_{6})\mid~=~\mid\mathscr{L}(D_{6}(a_{2}))\mid~=~\mid\mathscr{L}(D_{6}(a_{1}))\mid~\leq~66+12=78.\\$
Let the coordinates of roots of $E_{7}$ be as follows
[…]
Consider maximal and minimal roots $\pm\mu$ in $E_{7}$. Here, $\pm\mu\in\varPhi(E_{7})\backslash\varPhi(D_{6})$ and $\pm\mu_{max}\not\in\varPhi(D_{6})$. Note that $\pm\mu$ are orthogonal to any simple root except for $\tau_{6}$. Then vectors $\pm\mu^{\nabla}$ are zero, i.e.,
$\mid\mathscr{L}(D_{6})\mid~=~\mid\mathscr{L}(D_{6}(a_{1}))\mid~=~\mid\mathscr{L}(D_{6}(a_{2}))\mid~\leq~78-2=76.\\$ (5.8)
The full linkage system for $\mathscr{L}(D_{6})$ consists of the $12$-element linkage system $\mathscr{L}_{D_{7}}(D_{6})$ (Lemma 5.1(iii)) and two orbits of the $E$-component, each of which contains $32$ linkages. These $E$-components are depicted in Fig. 12, they contain $2\times 32=64$ elements. Thus, $D$-component and $E$-components together contain $12+64=76$ elements in $\mathscr{L}(D_{6})$. So, in (5.8) all full linkage systems $\mathscr{L}(D_{6})$, $\mathscr{L}(D_{6}(a_{1}))$ and $\mathscr{L}(D_{6}(a_{1}))$ contain exactly $76$ elements.
(v) For $l=7$, we have
$\mid\mathscr{L}_{E_{8}}(D_{7}(a_{2}))\mid~=~\mid\mathscr{L}_{E_{8}}(D_{7}(a_{1}))\mid~=~\mid\mathscr{L}_{E_{8}}(D_{7})\mid~\leq~\mid\varPhi(E_{8})\mid-\mid\varPhi(D_{7})\mid~\leq~240-84=156.\\$
By Lemma 5.2(i), in $\mathscr{L}_{E_{8}}(D_{7})$ there are $14$ pairs having the same linkage label vectors, then
$\mid\mathscr{L}_{E_{8}}(D_{7}(a_{2}))\mid~=~\mid\mathscr{L}_{E_{8}}(D_{7}(a_{1}))\mid~=~\mid\mathscr{L}_{E_{8}}(D_{7})\mid~\leq~156-14=142.\\$
In fact, $\mid\mathscr{L}_{E_{8}}(D_{7})\mid$ contains only $128$ elements in two $64$-element orbits which can be found in [St14, Figs. C.63, C.64]. The full linkage system for $\mathscr{L}(D_{7})$ consists of the $14$-element linkage system $\mathscr{L}_{D_{8}}(D_{7})$ (Lemma 5.1(iii)) and two orbits of the $E$-component. The full linkage systems $\mathscr{L}(D_{7})$, $\mathscr{L}(D_{7}(a_{1}))$ and $\mathscr{L}(D_{7}(a_{2}))$ contain exactly $142$ elements. ∎
Figure 9. For $l>4$, the partial linkage system $\mathscr{L}_{D_{l+1}}(D_{l})$ contains $2l$ linkages. For $l>7$, $\mathscr{L}_{D_{l+1}}(D_{l})$ coincides with the full linkage system $\mathscr{L}(D_{l})$.
Figure 10. For $l>4$, the partial linkage system $\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))$ contains $2l$ linkages. For $l>7$, $\mathscr{L}(D_{l}(a_{k}))=\mathscr{L}_{D_{l+1}}(D_{l}(a_{k}))$ .
Figure 11. Two orbits of the $E$-component of the linkage system $\mathscr{L}(D_{5})$. Each orbit consists of two loctets.
Figure 12. Two orbits of the $E$-component of the linkage system $\mathscr{L}(D_{6})$. Each orbit consists of four loctets.
Appendix A Gabrielov transformations and the Ovsienko theorem
A.1. Gabrielov transformations
In 1973, A. Gabrielov introduced mappings changing quadratic forms associated with singularities, [G73, §6]. Gabrielov’s example corresponding to the singularity $x^{3}+y^{3}+z^{2}$ is shown in Fig. 13. Essentially, Gabrielov transformations in Fig. 13 similar to transitions of §1.6. The sequence of transformation $T^{(i)},i=1,2,3$ modify the underlying diagram to the Dynkin diagram $D_{4}$, change the basis (vanishing cycles) and the matrix of the quadratic form (intersection matrix) which is similar to a certain Cartan matrix, [E18], [Go99]. Gabrielov transformations are used for classification of some positive definite forms in [DD95].
Figure 13. Gabrielov’s example of changing the basis corresponding to a singularity $x^{3}+y^{3}+z^{2}$. Transitions $T^{(i)},i=1,2,3$ act only on one basis vector and fix all others.
Figure 14. The Gabrielov example of changing the basis corresponding to a singularity $x^{4}+y^{3}+z^{2}$. Transitions $T^{(i)},1\leq i\leq 7$ act only on one basis vector and fix all others.
Another example corresponds to the singularity $x^{4}+y^{3}+z^{2}$, see Fig. 14. The sequence of Gabrielov transformations $T^{(i)},1\leq i\leq 7$ in Fig. 14 modify the underlying diagram to the Dynkin diagram $E_{6}$.
Let $B=B_{\Gamma}$ be the (partial) Cartan matrix associated with the diagram $\Gamma$, $E_{ij}$ be the elementary $n\times n$ matrix with unique nonzero entry a $1$ at slot $(i,j)$, and $G^{B}_{ij}$ - the linear transformation given by matrix $G^{B}_{ij}=I-b_{ij}E_{ij}$. For each $i\neq j$ in $\{1,\dots,n\}$, the Gabrielov transformation $\mathscr{G}_{ij}$ is defined as the following composition:
$B\longrightarrow\mathscr{G}_{ij}(B)=BG^{B}_{ij},$ (A.1)
see [BGP19, Prop. 2.17].
A.2. Inflations and Ovsienko’s theorem
Let $\varepsilon$ be a sign $\varepsilon\in\{+,-\}$, and for $1\leq i,j\leq n$ define the linear transformation $T^{\varepsilon_{ij}}:\mathbb{Z}^{n}\longmapsto\mathbb{Z}^{n}$ by
$T^{\varepsilon}_{ij}:v\longmapsto v-\varepsilon v_{i}\alpha_{j}.$
Note that $T^{+}_{ij}$ is the inverse of $T^{-}_{ij}$. The transformation $T^{-}_{ij}$ is called a deflation for $B_{\Gamma^{\prime}}$ if $b_{ij}<0$, and $T^{+}_{ij}$ is called an inflation for $B_{\Gamma^{\prime}}$ if $b_{ij}>0$. Inflations and deflations are called flations. A finite composition of flations is an iterated flation:
$T=T^{\varepsilon_{1}}_{i_{1}j_{1}}\dots T^{\varepsilon_{r}}_{i_{r}j_{r}}.$ (A.2)
Composition (A.2) is inductively defined, see [BGP19, §2.4].
The inflations technique introduced by S. Ovsienko in [O78] in the context of weakly positive unit quadratic forms, see [K05, Rem. 4.4] and [MM19, Exam. 5.2], is similar to Gabrielov’s transformations. If $\lvert b_{ij}\rvert=1$ and $\varepsilon\in\{+,-\}$ is such that $\lvert b_{ij}\rvert=\varepsilon b_{ij}$ then the quadratic form $BT^{\varepsilon}_{ij}$ coincides with the Gabrielov transformation $\mathscr{G}_{ij}(B)$ from (A.1).
Theorem (Ovsienko).
Let $B$ be a positive definite unit form. Then there exists an iterated inflation $T$ and a unique (up to permutation of components) disjoint union of Dynkin diagrams $\Gamma$ such that $BT$ coincides with the Cartan matrix $B_{\Gamma}$.
For the proof, see [BGP19, Th. 2.20]. For the further development of the inflation mappings, see [BP99], [H98], [K05]. [MM19], [MSZ17].